OpenStudy (anonymous):

REALLY SIMPLE FAST QUESTION + MEDAL 4 ANSWER

OpenStudy (anonymous):

How do you square this lmao: \[\sqrt{4h ^{2}+4}\]

OpenStudy (anonymous):

like if i were to square that would i use the foil method or would it just be \[4h ^{2} + 4\]

OpenStudy (anonymous):

@DanJS @Hero @ParthKohli

OpenStudy (caozeyuan):

4h^2+4 is correct

OpenStudy (anonymous):

Yeap, since \[(\sqrt(3))^2=3\]